# GTPO: Trajectory-Based Policy Optimization in Large Language Models

### v3

@@ 6.2 Ablation Studies
Entropy-based terms. To evaluate the impact of entropy-based terms, Figure 4 shows the training curves for average completion entropy (left), accuracy (middle), and formatting (right) for LLaMA-8B on the MATH dataset. We compare different entropy regularization strengths: $\gamma=0.1$, $\gamma=0.01$, $\gamma=0.001$, $\gamma=10^{-6}$, and $\gamma=0.1$ without applying the filtering defined in Eq. 6 (denoted as “NO $\delta_{i}$”). This last setting highlights the impact of not filtering out high-entropy completions that could trigger policy collapse.
As shown in the figure, larger values of $\gamma$ lead to improved accuracy and formatting, with $\gamma=0.1$ achieving the highest accuracy overall. In contrast, when filtering is disabled ($\gamma=0.1$, No $\delta_{i}$), both accuracy and formatting collapse, highlighting the critical role of the filtering term in maintaining training stability. This behavior is further supported by the entropy plot (left side): without filtering, entropy steadily increases and remains above $\ln 2$, eventually destabilizing formatting (initially, as seen in the red region) and later affecting accuracy. In contrast, when filtering is applied, entropy remains below $\ln 2$ and gradually decreases over time.
Interestingly, the figure shows that higher values of $\gamma$ lead to both greater stabilized entropy (left plot) and improved performance. Regarding performance, in terms of accuracy and formatting, this trend can be explained by the fact that very low entropy causes the model to become overly confident, limiting its ability to explore (e.g., with $\gamma=0.001$, where entropy continues to decrease and accuracy plateaus around step 4000). In contrast, having moderate entropy promotes continued exploration during training, resulting in more diverse and informative completions.
To better understand the behaviour of the entropy curves, where larger $\gamma$ values appear to converge toward higher entropy values, when a completion receives a negative advantage, entropy tends to increase (see Eq. 4). In this context, a stronger entropy regularization term (Eq. 7) amplifies the negative gradient on the selected tokens, increasing the probability of the unselected ones. This flattens the output distribution slightly, encouraging the model to continue exploring even in the later training stages.
@@
Conflict-Aware gradient correction. Figure 5 shows the accuracy and formatting training curves of LLaMA on GSM8K. The model is trained using GRPO with KL $\beta$ values set to 0, 0.04 (Shao et al. 2024), and $10^{-6}$, while for GTPO, we use the full version (Eq. 7) and a variant without entropy-based filtering and regularization (denoted as “No $\langle H\rangle_{i}$ - No $\delta_{i}$" in the figure). This latter configuration isolates the effect of GTPO when relying solely on the Conflict-Aware Gradient Correction component (i.e., Eq. 4). As shown in the figure, GTPO outperforms GRPO in both accuracy and formatting during the first 2,500 steps. Beyond this point, GTPO with regularization and filtering continues to maintain better performance, whereas the variant without these components begins to degrade and eventually falls below GRPO. This behavior is expected, as the absence of regularization prevents the model from balancing the impact reward signals over time. Most importantly, before the policy collapse, the use of gradient correction alone yields higher rewards than GRPO, highlighting its benefits.
@@ 7 Conclusion
In this work, we presented GTPO (Group-relative Trajectory-based Policy Optimization), a stable and effective policy optimization method for language models. GTPO addresses two key issues identified in GRPO: gradient conflicts affecting shared tokens in group of completions, and policy collapse. The core idea of GTPO is to control the divergence of completions within the same group, considering them as linked trajectories. This is achieved by mitigating gradient issues through the identification and masking of conflict tokens, and addressing policy collapse via entropy-based filtering and regularization. Through comprehensive experiments on GSM8K, MATH, and AIME2024, we demonstrated that GTPO consistently outperforms both GRPO and SFT across multiple settings.
A promising future direction will be to further investigate a theoretical minimum entropy threshold, with additional insights provided in the appendix, which may guide models toward an optimal entropy level and exploration. We believe that the insights presented in this study offer important contributions into the understanding of stable model alignment within the learning dynamics of language models, particularly in relation to entropy bounds and gradient conflicts that can arise in group-relative policy optimization paradigms.
Appendix and supplementary material for the paper :
GTPO: Trajectory-Based Policy Optimization in Large Language Models
8 Extended analysis of GRPO and GTPO
8.1 The Role of KL Divergence on GRPO
We analyze the impact of the KL divergence term on the aggregated GRPO loss by transitioning from a token-level formulation to a trajectory-level one. Recall the normalized advantage function used in GRPO:
@@ 8.1 The Role of KL Divergence on GRPO
$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{R_{i}-\bar{R}}{\operatorname{std}(R)}-\beta\cdot\mathrm{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)$ (9)
@@
$\tilde{\mathcal{J}}_{\mathrm{GRPO}}(\theta):=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{R_{i}-\bar{R}}{\operatorname{std}(R)}.$ (10)
@@
$\tilde{\mathcal{J}}_{\mathrm{GRPO}}(\theta)$ $=\frac{1}{G}\sum_{i=1}^{G}\left(\frac{R_{i}}{\operatorname{std}(R)}-\frac{\bar{R}}{\operatorname{std}(R)}\right)$
$=\frac{1}{\operatorname{std}(R)}\left(\frac{1}{G}\sum_{i=1}^{G}R_{i}-\bar{R}\right)$
$=\frac{1}{\operatorname{std}(R)}(\bar{R}-\bar{R})=0.$
@@
$\mathcal{J}_{\text{GRPO}}(\theta)=-\beta\cdot\mathrm{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right).$ (11)
@@
8.2 GRPO Gradient
Given the GRPO objective defined in Equation 1, we aim to calculate the gradient. Similar to what we have done in GRPO Loss, we decompose the objective by focusing our attention on Eq. 10.
$\tilde{\mathcal{J}}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$ (12)
In this case, to simplify the calculation, $\hat{\mathcal{A}}_{i}$ is kept unresolved (Eq. 8), and the term $\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$ is included to ensure the correct computation of the gradient. However, as shown in Section 3, this term is equal to 1 for every $t\in|o_{i}|$ because only a single iteration is considered, implying that $\pi_{\theta}=\pi_{\theta_{\text{old}}}$.
$\nabla_{\theta}\tilde{\mathcal{J}}_{\text{GRPO}}(\theta)$ $=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\nabla_{\theta}\left[\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\right]$
$=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{\nabla_{\theta}\left[\pi_{\theta}(o_{i,t}|q,o_{i,<t})\right]}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}$
@@ 8.2 GRPO Gradient
$\nabla_{\theta}\tilde{\mathcal{J}}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\cdot\\ \cdot\nabla_{\theta}\left[\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t}))\right]$ (13)
Then considering that the ratio between $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ is equal to 1, we can rewrite the complete gradient as:
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)$ $=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\nabla_{\theta}\left[\log(\pi_{\theta}(o_{i,t}|q,o_{i,<t}))\right]$
$-\beta\cdot\nabla_{\theta}\left[\mathrm{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right]$ (14)
How gradients affect distribution over tokens.
LLMs typically terminate with a softmax layer, which produces a probability distribution over the vocabulary. To simplify notation in what follows, we denote $(o_{i,t}|q,o_{i,<t})$ as $(t^{\prime})$, yielding the following expression for the output distribution:
@@
Substituting this into the GRPO gradient (ignoring the KL term for clarity) yields:
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}$ $(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Bigg{(}\nabla_{\theta}f^{k}_{\theta}(1-\pi_{\theta}(t^{\prime}))$
@@
$\left.-\sum_{k\neq t}^{V}\pi_{\theta}(k^{\prime})\nabla_{\theta}f^{k}_{\theta}\right)-\beta\cdot\nabla_{\theta}\left[\mathrm{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right]$ (19)
8.3 Impact of KL $\beta$-coefficient on GRPO

### v4

@@ 6.2 Ablation Studies
Conflict-Aware gradient correction. Figure 3 shows the accuracy and formatting training curves of LLaMA on GSM8K. The model is trained using GRPO with KL $\beta$ values set to 0, 0.04 GRPO, and $10^{-6}$, while for GTPO, we use the full version (Eq. 11) and a variant without entropy-based filtering and regularization (denoted as “No $\langle H\rangle_{i}$ – No $\delta_{i}$ in the figure). This latter configuration isolates the effect of GTPO when relying solely on the Conflict-Aware Gradient Correction component (i.e., Eq. 8). As shown in the figure, GTPO outperforms GRPO in both accuracy and formatting during the first 2,500 steps. Beyond this point, GTPO with regularization and filtering continues to maintain better performance, whereas the variant without these components begins to degrade and eventually falls below GRPO. This behavior is expected, as the absence of regularization prevents the model from balancing the impact reward signals over time. Most importantly, before the policy collapse, the use of gradient correction alone yields higher rewards than GRPO.
Impact of Entropy Regularization. Figure 4(a) shows the training curves for average completion entropy (left), accuracy (middle), and formatting (right) on LLaMA 8B trained on MATH. We compare several entropy regularization strengths ($\gamma=0.1,0.01,0.001,10^{-6}$) and a variant with $\gamma=0.1$ but without entropy filtering (“NO $\delta_{i}$”). Higher values of $\gamma$ generally yield better accuracy and formatting, with $\gamma=0.1$ performing best. When filtering is removed, however, both metrics collapse despite using the same $\gamma$, confirming that filtering high-entropy completions is essential for stability. The entropy curves further clarify this behavior. Without filtering, entropy continually rises above $\ln 2$, destabilizing formatting first and accuracy later. With filtering, entropy stays below $\ln 2$ and decreases smoothly. Moderate entropy (achieved with larger $\gamma$) supports exploration, producing more diverse and informative completions; excessively low entropy (e.g., with $\gamma=0.001$) leads to overconfidence and early performance saturation. Finally, note that larger $\gamma$ values tend to maintain slightly higher entropy. When a completion receives a negative advantage, entropy naturally increases (Eq. 8), and a stronger regularization term (Eq. 11) amplifies this effect by slightly flattening the token distribution. However, this increase remains controlled and does not destabilize training; instead, it preserves a small amount of exploration that remains beneficial even in later stages.
$\langle H\rangle_{\text{i}}$ Threshold in GTPO. Figure 4(b) analyzes the effect of different entropy thresholds $\langle H\rangle_{\text{i}}$ on GTPO training for LLaMA on GSM8K. We consider three filtering settings: $0.7$, $1.05$, and $2.0$, which roughly correspond to output distributions over, respectively, 2, 3, and 4 equiprobable tokens (with no residual probability mass). With the stricter thresholds (0.7 and 1.05), the entropy plot shows that the average completion entropy remains low and stable, while the loose threshold (2.0) lets many high-entropy completions pass, causing entropy to drift toward the instability region around $\ln 2$ and destabilizing the policy. The accuracy curves mirror this behavior: tighter thresholds yield smoother improvements and higher final accuracy, whereas the loose threshold leads to stagnation. The formatting plot further confirms that insufficient filtering (as in the 2.0 setting) quickly degrades structural quality, whereas stricter thresholds preserve formatting and thus support both stability and output quality.
Effects of Negative-Only Entropy Filtering. Figure 4(c) compares GRPO against GTPO variants using Qwen on GSM8K dataset, analyzing different entropy filtering strategies. In OnlyNeg setting, the entropy threshold, tested at $\langle H\rangle_{\text{i}}<2.00$, <1.05<1.05 and <0.7<0.7, is applied only to completions with negative advantages, meaning all positive completions are retained. Instead the standard approach for Qwen (pink curve) does not apply the completion filtering. The results show that restricting filtering only to negative completions worsens performance; it drives entropy too low, reducing exploration, while failing to filter high entropy positive completions that may reinforce erroneous trajectories.
Effect of Regularization and Filtering on GRPO. Figure 4(d) compares LLaMA on GSM8K under three settings: (i) standard GRPO without KL ($\beta=0$), (ii) GRPO with entropy filtering ($\langle H\rangle_{i}<1.05$) and entropy regularization ($\gamma=0.1$), and (iii) GTPO with the same entropy constraints. The entropy plot (left) shows that standard GRPO without KL is highly unstable: entropy quickly spikes and then collapses, leading to a complete loss of structure. This is also reflected in the performance curves, where accuracy remains low and formatting breaks down early. Adding entropy regularization and filtering (roughly corresponding to three equiprobable tokens) significantly improves the situation. The entropy stays within a stable range, accuracy rises to reasonable levels, and formatting recovers. This shows that entropy control alone does not fully fix GRPO, but it substantially improves its stability and prevents the immediate collapse observed in the baseline. However, GTPO performs consistently better even under the same constraints. It maintains a lower and more stable entropy, achieves higher final accuracy, and preserves near-perfect formatting throughout training. In other words, while entropy control stabilizes GRPO, GTPO provides an additional layer of robustness and leads to the best overall performance.
@@ 7 Conclusion
In this work, we introduced Group-relative Trajectory-based Policy Optimization (GTPO), which addresses two issues of GRPO, gradient conflicts and policy collapse, through conflict-aware updates and entropy control. We evaluated GTPO using LLaMA 8B and Qwen 2.5 (3B) on mathematical reasoning tasks, covering in-distribution datasets (GSM8K, MATH) and out-of-distribution benchmarks (AIME 2024, AIME 2025, AMC 2023). Through extensive ablation studies, we analyzed hyperparameter sensitivity to show the impact of each GTPO Objective component and how some of these components (entropy regularization and completion filtering) can also improve the performances of GRPO. On in-ditribution datasets, our results demonstrate that GTPO significantly outperforms baselines, improving Pass@32 and Maj@32 by up to 10 percentage points over GRPO and SFT. On out-of-distribution benchmarks, the gains are even larger, up to 75 points on Pass@64 and 37.5 points on Maj@64 over GRPO, and 32.5 and 42.5 points over SFT. For future work, we plan to further identify theoretical entropy bounds to ensure effective exploration without compromising stability.
Appendix
8 Proofs and Repricability
8.1 Group-Relative Policy Optimization
The Role of KL Divergence on GRPO. We analyze the impact of the KL divergence term on the aggregated GRPO loss by transitioning from a token-level formulation to a trajectory-level one, when $r_{i,t}=1$. Recall the normalized advantage function used in GRPO-like:
@@ 8.1 Group-Relative Policy Optimization
$\mathcal{J}_{\mathrm{GRPO}}(\theta)$ $=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{R_{i}-\bar{R}}{\operatorname{std}(R)}$
$-\beta\cdot\mathrm{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\Bigg]$ (13)
@@
$\tilde{\mathcal{J}}_{\mathrm{GRPO}}(\theta):=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{R_{i}-\bar{R}}{\operatorname{std}(R)}\Bigg]$ (14)
@@
$\tilde{\mathcal{J}}_{\mathrm{GRPO}}(\theta)$ $=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\left(\frac{R_{i}}{\operatorname{std}(R)}-\frac{\bar{R}}{\operatorname{std}(R)}\right)\Bigg]$ (15)
$=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{\operatorname{std}(R)}\left(\frac{1}{G}\sum_{i=1}^{G}R_{i}-\bar{R}\right)\Bigg]$
$=\mathbb{E}_{q,\{o_{i}\}}[0].$
@@
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q,\{o_{i}\}}\Big[-\beta\cdot\mathrm{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\Big]$ (16)
@@
GRPO Gradient. Given the GRPO objective defined in Equation 1, we aim to calculate the gradient.
$\tilde{\mathcal{J}}_{\text{GRPO}}(\theta)=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{\pi_{\theta}(o_{i,t}\mid s_{i,t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid s_{i,t})}\Bigg]$ (17)
To simplify the calculation, we keep $\hat{\mathcal{A}}_{i}$ in its unexpanded form (Eq. 12) and reduce the term $C_{i,t}$,(Eq. 2) to the ratio $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid s_{i,t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid s_{i,t})}$. This substitution relies on the fact that the objective function’s clipping mechanism eliminates gradient contributions from tokens where $|r_{i,t}(\theta)|>1+\epsilon$. Therefore, the summation is performed exclusively over the set of active (non clipped) tokens. Formalizing the excluded tokens adds notational burden with no impact on the resulting gradient analysis.
$\nabla_{\theta}$ $\tilde{\mathcal{J}}_{\text{GRPO}}(\theta)=$
$=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\nabla_{\theta}\left[\frac{\pi_{\theta}(o_{i,t}\mid s_{i,t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid s_{i,t})}\right]\Bigg]$
$=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\frac{\pi_{\theta}(o_{i,t}\mid s_{i,t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid s_{i,t})}\Bigg]$
@@
$\nabla_{\theta}\tilde{\mathcal{J}}_{\text{GRPO}}(\theta)=$
$\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}r_{i,t}(\theta)\cdot\nabla_{\theta}\left[\log(\pi_{\theta}(o_{i,t}|s_{i,t}))\right]\Bigg]$ (18)
We can rewrite the complete gradient as:
$\nabla_{\theta}\mathcal{J}_{\text{GRPO}}(\theta)=$
$\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}r_{i,t}(\theta)\nabla_{\theta}\left[\log(\pi_{\theta}(o_{i,t}|s_{i,t}))\right]$
$-\beta\cdot\nabla_{\theta}\left[\mathrm{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right]\Bigg]$ (19)
How gradients affect distribution over tokens. LLMs typically terminate with a softmax layer, which produces a probability distribution over the vocabulary. To simplify notation in what follows, we denote $(o_{i,t}|q,o_{i,<t})$ as $(t^{\prime})$, yielding the following expression for the output distribution:
@@
Substituting this into the GRPO gradient yields:
$\nabla_{\theta}$ $\mathcal{J}_{\text{GRPO}}(\theta)=$
$\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{\hat{\mathcal{A}}_{i}}{|o_{i}|}\sum_{t=1}^{|o_{i}|}r_{i,t}\Bigg(\nabla_{\theta}f^{k}_{\theta}(1-\pi_{\theta}(t^{\prime}))$
@@
$\left.-\sum_{k\neq t}^{V}\pi_{\theta}(k^{\prime})\nabla_{\theta}f^{k}_{\theta}\right)-\beta\cdot\nabla_{\theta}\left[\mathrm{D}_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\mathrm{ref}}\right)\right]\Bigg]$ (24)
8.2 Group-relative Trajectory-aware Policy Optimization: Objective
This section formalizes the Conflict-Aware Gradient Correction for GTPO, deriving the token-level loss and gradient. GTPO refines GRPO by weighting each token with a coefficient $\lambda_{i,t}$, determined by conflict regions and the advantage $A_{i}$. Furthermore, to mitigate instability, we incorporate an entropy-based filter $\delta_{i}$. The resulting objective function is defined as:
$\mathcal{J}^{*}:=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t\in o_{i}}r_{i,t}\lambda_{i,t}A_{i}\Bigg]$
$=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\left(\sum_{t\in c_{i}}r_{i,t}\lambda_{i,t}A_{i}+\sum_{t\in u_{i}}r_{i,t}\lambda_{i,t}A_{i}\right)\Bigg]$
Where $o_{i}$ denotes the set of tokens in each completion, $c_{i}$ and $u_{i}$ represent the subsets of conflicting and non-conflicting tokens, respectively. The notation $|\cdot|$ indicates the cardinality (number of elements) of each set and $\lambda_{i,t}$ is a parameter designed to account for conflict and non-conflict formulation. To motivate the definition of $\lambda_{i,t}$, we assume the experimental condition where $r_{i,t}=1$. Furthermore, to ensure that the model retains the characteristics of GRPO for non-conflicting tokens, we define $\lambda_{i,t}$ as follows:
$\mathcal{J}^{*}=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\left(|c_{i}|\lambda_{i}A_{i}+\left(|o_{i}|-|c_{i}|\right)A_{i}\right)\Bigg]$
$=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\left(\sum_{i=1}^{G}\frac{|c_{i}|}{|o_{i}|}\lambda_{i}A_{i}+\sum_{i=1}^{G}A_{i}-\sum_{i=1}^{G}\frac{|c_{i}|}{|o_{i}|}A_{i}\right)\Bigg]$
$=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\left(\sum_{i=1}^{G}\frac{|c_{i}|}{|o_{i}|}A_{i}(\lambda_{i}-1)+\sum_{i=1}^{G}A_{i}\right)\Bigg]$
$\mathcal{J}^{*}=\mathbb{E}_{q,\{o_{i}\}}\Bigg[$ $\frac{1}{G}\Bigg(\sum_{i=1}^{G^{+}}\frac{|c_{i}|}{|o_{i}|}A_{i}(\lambda_{i}-1)$
$+\sum_{i=1}^{G^{-}}\frac{|c_{i}|}{|o_{i}|}A_{i}(\lambda_{i}-1)\Bigg)\Bigg]$
$\mathcal{J}^{*}$ $=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\left(\sum_{i=1}^{G^{+}}\frac{|c_{i}|}{|o_{i}|}|A_{i}|(\lambda-1)+\sum_{i=1}^{G^{-}}\frac{|c_{i}|}{|o_{i}|}|A_{i}|\right)\Bigg]$
The resulting equation for the aggregated loss is:
$\mathcal{J}^{*}$ $=\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{|c_{i}|}{|o_{i}|}|A_{i}|\Bigg]$ (25)
Instead the resulting equation for the token level loss, considering $r_{i,t}\neq 1$, is:
$\mathcal{J}^{*}=$ $\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=0}^{o_{i}}r_{i,t}\lambda_{i,t}A_{i}\Bigg]$ (26)
where:
[…]
GTPO Gradient: To calculate the gradient of our proposed loss, we should start from equation 26, then we will apply the same tricks used to calculate GRPO loss (eq. 17).
$\nabla_{\theta}\mathcal{J}^{*}=$ $\;\mathbb{E}_{q,\{o_{i}\}}\Bigg[\frac{1}{G}\sum_{i=1}^{G^{+}}\frac{A_{i}}{|o_{i}|}\left(2\sum_{t\in c_{i}}g_{i,t}+\sum_{t\in u_{i}}g_{i,t}\right)$
$+\frac{1}{G}\sum_{i=1}^{G^{-}}\frac{A_{i}}{|o_{i}|}\sum_{t\in u_{i}}g_{i,t}\Bigg]$ (28)
where:
$g_{i,t}=r_{i,t}\nabla_{\theta}\left[\log\left(\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})\right)\right]$ (29)
Reward Settings. The reward used in this paper is composed of two components: the first one refers to the format of the model’s answer, and the second one refers to the accuracy of the model’s answer. The first component assesses formatting: the model receives a score of 10 if it includes all predefined special formatting tokens specified in the prompt (see box 8.2). If only a subset of the reasoning or answer tokens are present, the model is awarded 1 point; otherwise, it scores 0. The second component evaluates accuracy: the model obtains a score of 10 if the correct target answer is correctly enclosed within the answer tokens; otherwise, it receives 0. The final reward is the sum of these two components and is provided to guide the learning algorithms.
Training Settings. Listing 1 reports the main training parameters used for GRPO, GTPO and SFT, for both LLaMA and Qwen models. To implement GTPO we used Unsloth unsloth.
model_name: "meta-llama/meta-Llama-3.1-8B-Instruct" / "Qwen/Qwen2.5-3B-Instruct"
max_seq_length: 5500
max_prompt_length: 4000
lora_rank: 128 / 64 (Qwen)
load_in_4bit: true
gpu_memory_utilization: 0.4
target_modules:
- "q_proj"
- "k_proj"
- "v_proj"
- "o_proj"
- "gate_proj"
- "up_proj"
- "down_proj"
random_seed: 3407
warmup_ratio: 0.005
learning_rate: 1e-6
adam_beta1: 0.999999 / 0.9 (GRPO & SFT)
adam_beta2: 0.999999 / 0.95 (GRPO & SFT)
weight_decay: 0.1
beta: 0.0/0.04/10e-6 (GRPO) / "None" (GTPO & SFT)
lr_scheduler_type: "cosine"
optimizer: "paged_adamw_8bit"
logging_steps: 1
per_device_train_batch_size: 1
gradient_accumulation_steps: 1
num_generations: "8/12" / "None" (SFT)
num_train_epochs: 1
num_iterations: 1
save_steps: 500
max_grad_norm: 0.1
report_to: ["wandb"]
Listing 1: Training settings for GTPO, GRPO and SFT.
9 Complementary results
9.1 Impact of KL $\beta$-coefficient on GRPO.